What is the slope of the line:
4x-5y=30?

Answer by tutorcecilia(2152)

4x-5y=30
4x-4x-5y=-4x+30
-5y=-4x+30
-5y/-5=-4x/-5+30/-5
y=(4/5)x-6
Slope=m=4/5